# Neural Networks for Fast Optimisation in Model Predictive Control: A Review

### v2

@@ 6.1 Methods without guarantees
As for LMPC and NMPC, the simplest way to approximate RMPC with an NN, is to use imitation learning to learn the behaviour of the controller to a satisfactory level of performance. Lucia and Karg Lucia and Karg (2018) demonstrated that this approach is applicable to RMPC despite its higher complexity. Specifically, they used a DNN with $tanh$ activation functions to learn the behaviour of multi-stage RMPC (msRMPC). Through simulation examples they proved the success of the approach at approximating the equivalent robust nonlinear controller for an industrial polymerisation reactor. The approximate controller was 2,800 times faster than the exact controller and could successfully be executed at the required rate in real time, however, it should also be pointed out that the average constraint violation, despite remaining low, was one order of magnitude larger than that of the exact controller. Thus, the pitfall of the basic imitation learning setup remains the same as previously discussed: it is not suitable for use in safety-critical systems because it provides no guarantees that the resulting approximate controller will respect the constraints of the system.
@@
In follow up work, Karg and Lucia Karg and Lucia (2019) presented a modified approach where instead of directly learning the behaviour of msRMPC, an NN was used to learn the behaviour of a slightly modified controller which contained a constraint back-off parameter that made the controller more conservative. Through simulation experiments on a small nonlinear plant with one input and three states, the authors showed that the modified imitation learning approach could be successfully used to create an approximate controller that respected the constraint to which the back-off was applied. In their example, the back-off parameter was applied to a scalar nonlinear constraint, no examples were provided for the case of multiple constraints on inputs and states. The revised method with back-off parameters is a reasonable compromise between controller conservativeness and safety. This is important because, even in systems that are not safety-critical, avoiding constraint violations may be desirable to prevent unexpected stops or mechanical failure. Nonetheless, it should be noted that the back-off strategy does not provide any formal guarantees of constraint satisfaction. By using it, constraint violation is less likely to occur and the authors outlined a method via which the probability of violation can be computed after the approximation is generated. Yet, it is important to note that the method itself does not inherently provide probabilistic or strict guarantees.
@@
Bonzanini et al. Bonzanini et al. (2021) further embraced the idea that the lack of strict safety guarantees can be overlooked in favour of other valuable compromises such as enabling the use of real-time msRMPC controllers that use augmented models equipped with Gaussian Processes (GP) to learn and mitigate the effects of plant-model mismatch. The incentive for creating and using these augmented models is that the standard formulation of msRMPC assumes that the structure and distribution of uncertainties is bounded and known in advance. However, as the authors pointed out, uncertainty can be time-varying and some systems are just too hard to model using conventional approaches. In such cases GP regression can be used to learn the unmodeled dynamics or time-varying uncertainties of the system from data. The msRMPC formulation can then be modified to accommodate the GP-augmented models. Via simulation experiments on a system with significant plant-model mismatch, Bonzanini et al. showed that their learning-based reformulation of msRMPC outperformed the standard formulation but its computation cost remained too high for use in real time. To overcome this issue, they applied the basic imitation learning solution and trained a DNN to learn the behaviour of the reformulated controller. The resulting approximation enjoyed no safety guarantees, but its performance on closed-loop simulation examples matched that of the exact reformulated msRMPC controller at a quarter of the online computation cost. Karg and Lucia Karg and Lucia (2021) proposed a different method that achieves a similar result as Bonzanini et al. Bonzanini et al. (2021), instead of reformulating msRMPC to account for more complex models, they trained a DNN via imitation learning to approximate the standard form of msRMPC and then used the resulting network as an initialisation of RL. Further training of the initialised agent then took care of the problem of learning the unmodeled dynamics and time-varying uncertainties. Both of these studies Bonzanini et al. (2021); Karg and Lucia (2021), are examples of accepting the lack of guarantees of the simplest NNBO approach to enable the use of complex controllers that build beyond classical MPC formulations.
@@ 6.2 Methods with probabilistic guarantees
In this section, we present two frameworks that are designed to preserve certain theoretical properties of RMPC on a probabilistic basis. The first framework, proposed by Hertneck et al. Hertneck et al. (2018), is an extension of the pure regression approach which includes a validation step that once passed, equips the basic approximation with statistical guarantees on constraint satisfaction and stability. In short, the validation step tests a candidate DNN approximation of the original controller on several closed-loop trajectories which start and end in feasible regions of the system’s state space, and which the original RMPC controller can solve without violating any constraints. For each trajectory in the test set, an indicator function is applied which outputs 1 if the closed-loop solution of the DNN is within a specified tolerance of the solution of the real controller, or 0 if it is not. The mean of the indicator outputs is equivalent to $\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ in HI (6). It follows that, if the value of $E(\xi)$ and the RHS of (6) are predefined, then one can calculate $\epsilon$ and verify if the difference between $\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ and $\epsilon$ is greater or equal than the prescribed $E(\xi)$ for the current number of test trajectories. If the verification is not successful, then the number of test trajectories can be increased or the quality of the DNN approximation can be improved by modifying the DNN and re-training until the validation step succeeds. If a candidate solution passes the validation step, it carries the guarantee that with some predefined confidence, its solutions are stable and constraint compliant (as defined by the indicator function) with a probability of at least the selected $E(\xi)$.
@@
Hertneck et al. Hertneck et al. (2018) applied this methodology to an RMPC formulation designed for robustness against bounded input disturbances. To preserve said robustness on the approximate controller, their choice of indicator function outputs 1 if the infinity norm of the input error between the real and approximate controllers over a closed loop trajectory is less or equal to the input disturbance bound of the original controller. Otherwise the indicator outputs 0. The effectiveness of the approach was demonstrated on a simulation study entailing the control of a continuously stirred tank reactor, in said study the performance of the approximate controller was found to be indistinguishable from that of the original policy while being over 200 times faster to evaluate.
@@
Later on, Nubert et al. Nubert et al. (2020) applied the same approach to a more complex version of RMPC designed for setpoint tracking of nonlinear systems with robustness against plant-model mismatch. In this case the choice of indicator function was adjusted to output 1 if the solutions of the approximate controller complied with the maximum admissible model mismatch for the approximation, and 0 otherwise. The effectiveness of the reformulated approach was demonstrated on setpoint tracking experiments with a real robot arm that was represented by a simplified model with 8 states, 4 inputs and 3 outputs for control purposes. Nubert et al. Nubert et al. (2020) reported difficulties in satisfying their HI validation criterion for high probability levels but the resulting approximate controller worked well in practice and could be executed 200 times faster than the original controller thereby allowing the use of this complex form of RMPC in real time. It should be noted that this is the only framework that has been used across multiple studies and proved to be scalable and adaptable to more complex controllers while retaining the same speed up factor. It is also one of the few frameworks that have been tested on experiments with real hardware instead of simulation.
@@
Karg et al. Karg et al. (2021) proposed a different framework for obtaining statistical guarantees on approximations of msRMPC which is suitable for the case where the performance of a controller with respect to a scalar indicator depends on a set of hyper-parameters. In this framework a family of controller approximations is considered, where each member of the family corresponds to a controller with a particular combination of hyper parameters. Using concepts from order statistics and probabilistic robustness analysis, the authors came up with expressions from which the size of a test set can be computed as a function of the size of the family. For said test set the authors also derived expressions to calculate the number of tests that can be failed, the desired probability of indicator violation, and the desired confidence level of the probability of violation. Each controller in the finite family can be evaluated on this test set and if the number of failed tests is lower than the allowable number of failures, then the family member is guaranteed to carry the probabilistic properties that were prescribed. Karg et al. Karg et al. (2021) then showed how this framework could be applied to a family of DNN approximations where each family member was trained to approximate the behaviour of an msRMPC controller with a particular constraint tightening parameter. In this example, the scalar indicator function was set to the maximum constraint violation over a 60 $s$ simulation of the closed loop system and the target performance of the probabilistic validation was to allow no violation at all with a prescribed probability. By applying their framework the size of the test set was computed and the authors showed how the different members of the family met or failed the maximum number of failures requirement. A perk of this method is that it can be applied to any arbitrary controller, not just DNN approximations of msRMPC. However, for generating neural approximations of MPC, which is the focal point of this paper, the method requires creating a data set per family member being considered. That means evaluating expensive MPC policies on large data sets multiple times instead of once like in other approximation frameworks. Therefore, there seems to be no compelling reason to choose this approximation framework over the previously discussed method Hertneck et al. (2018); Nubert et al. (2020) given that both can yield the same probabilistic guarantees.
@@ 6.3 Methods with strict guarantees
In this section we present methods for approximating RMPC with DNN in manners that preserve some of the theoretical properties of MPC such as the guarantees of constraint satisfaction and stability. Daosud et al. Daosud et al. (2019) proposed an intuitive approach for preserving the stability and constraint satisfaction properties of msRMPC. Their method consists of training a DNN via imitation learning to predict the optimal cost of msRMPC for a given starting condition. They used this trained network as the terminal cost of a re-formulated msRMPC controller with a much shorter prediction horizon than the original. Shortening the prediction horizon greatly reduces the computational cost of optimisation but enables the re-formulated controller to retain the stability and constraint satisfaction properties of MPC because rigorous optimisation is still being conducted at each control interval. The authors demonstrated the effectiveness of the approach on a simulated semi-batch reactor problem consisting of a MIMO plant with dimensions as shown in Table 3. First, the original controller was implemented and simulated with a robust horizon of 1 step and a prediction horizon of 15 steps. Then, a DNN with $tanh$ activation functions was used to learn to predict the optimal cost of the solutions of this controller. Subsequently, the trained network was used as the terminal cost of a reformulated controller which had the same robust horizon as the original but a prediction horizon of only 4 steps. Through simulation experiments the authors verified that the new controller with the DNN terminal cost was stable, constraint compliant and ran 2.3 times faster on average than the original controller. The main advantages of this approach are its intuitive nature and that the computation time reductions should only improve as the length of the horizon of the original controller increases because the cost of evaluating the DNN would remain constant. The only disadvantage is that the optimality guarantees of MPC are not maintained. In their study Daosud et al. (2019), the authors reported a 5% performance loss when using the approximate controller compared to the original in terms of their control goal which was to maximise the concentration of a reactant in the semi-batch reactor. This performance loss was attributed to the prediction error of the DNN and it should be noted that the worst case or average performance loss were not reported, so in practice this number could be much higher, particularly if the quality of the DNN predictions worsens for harder problems and longer horizons. Nevertheless, while some performance is bound to be lost and the computation time reduction is not immense, the intuitiveness and guarantees of this method make it worthy of more investigation into its scalability and applicability to other types of MPC such as NMPC.
@@
The only other alternative with strict guarantees was proposed by Paulson et al. Paulson and Mesbah (2020) and relies on projection operations to preserve the constraint and stability guarantees of msRMPC. Their method consists of training a DNN to learn the mapping between states and control actions for standard msRMPC. Once the network is trained its outputs are projected via optimisation onto carefully selected sets that preserve certain properties of MPC. For preserving constraint satisfaction properties in the presence of uncertainties, the authors proposed projecting the DNN outputs onto the maximal RCI set. To preserve stability properties, the authors projected the results onto the maximal $\lambda$-contractive set, which is a set in state space where all states can be driven into a tighter region of the set by applying one control input in the absence of disturbances. With this approach one can have an approximate controller that is constraint compliant in the presence of disturbances and asymptotically stable in the absence of disturbances. Nonetheless, assuming an absence of disturbances defeats the purpose of using RMPC in the first place and while the use of the $\lambda$-contractive set can yield Input-to-State Stability (ISS) guarantees for disturbed cases, there are additional conditions that need to be satisfied in this case which are difficult and perhaps impossible to verify a priori. In fact, the requirements for the constraint satisfaction guarantees are also hard to meet because the RCI set cannot always be computed or approximated for complex system with large plants. The authors demonstrated their approach on a double integrator plant problem and reported a computation speed up of 400 times compared to the original controller when the optimisation based projection operation was solved offline as an mpQP. No speed up factors or results were reported for cases where the projection operation cannot be solved offline. The applicability of this method to more complex problems has not been studied but seems unlikely.
@@ 7.1 Problem types and the lack of benchmarks
From the graph it is clear that there are a few problem types that are significantly more popular, while other types have only been encountered once or twice. The trend we aim to highlight is the absence of well-established benchmarks for testing approximation frameworks. The stirred or semi-batch tank reactor problems are by far the most commonly encountered problem type in case studies. Ren et al. Ren et al. (2022) made the same observation on their survey on NN modelling approaches for MPC. Specifically, they noted that chemical reaction problems are commonly encountered in case studies of machine learning-assisted optimal control because it is difficult to model such processes from first principles. In the works covered in this review, versions of the stirred reactor problem type are usually polymerisation reactions or pH control problems. Although the plants involved vary in complexity, most are nonlinear MIMO plants of medium size with 4 to 8 states and 2 to 3 inputs. The second most popular problem type is the double integrator, which is a SISO plant with 2 states. This problem type is not dissimilar from undamped oscillator problems. We encountered some publications Paulson and Mesbah (2020); Karg and Lucia (2020b); Pin et al. (2010, 2013) that demonstrated their approach only on these problem types, which raises concerns about their scalability. Heating system problems are the third most popular type. The plants involved in these problems vary in size, ranging from medium to very large with hundreds of states. It is evident then that there is no standardised benchmark for showcasing the effectiveness of a given framework.
@@
Some research studies addressed the lack of benchmarks challenge by using multiple examples and more than one problem type to demonstrate an approach. For instance, Fabiani and Goulart Fabiani and Goulart (2022) used six versions of the oscillating masses problem with increasing dimensions and a fixed horizon length to demonstrate the scalability of their method. Similarly, Chen et al. Chen et al. (2022) considered three problem types and four case studies including the double integrator, quadrotor, and two instances of the oscillating masses problem. By demonstrating their approach on plants of different sizes and different nature, they not only proved the scalability of their method, but also its applicability beyond a specific problem type. Alternatively, Li et al. Li et al. (2022) focused on a single plant but made efforts to demonstrate the scalability of their method in different ways. Most notably, they showed how the training time and performance of their approach varied as more training data was considered. Studies such as these allow to make important comparisons between approximation methods. In turn, these comparisons shed light on the information that a comprehensive benchmark should aim to measure.
@@
Figure 5 shows several relationships which were estimated using the data extracted from the case studies in Fabiani and Goulart (2022); Chen et al. (2022); Li et al. (2022). Figures 5a – c compare the frameworks by Chen et al. Chen et al. (2022) and Fabiani and Goulart Fabiani and Goulart (2022) which are both of the strict guarantees type for LMPC, but are otherwise very different in nature. The former consists of training a DNN that predicts all primal variables to hot-start an optimisation solver. The latter proposes MILPs to compute the values of properties that can be used to verify the stability and constraint compliance of controllers approximated via DNN ReLU regression. As shown in the plots, these differences in nature have profound impacts on the way that both methods scale for problems of different size. For example, the equations of the fitted functions in Figure 5a show that the required network size in the method by Chen et al. Chen et al. (2022) scales exponentially with plant size. However, for the method by Fabiani and Goulart Fabiani and Goulart (2022) it scales linearly. This information is valuable because training large networks is likely to require more time and computational resources than for smaller networks. Note however, that because Fabiani and Goulart Fabiani and Goulart (2022) only considered one problem type of various sizes, it is not possible to say if this linear scaling holds in general. Figure 5b shows how the required NN size varies with prediction horizon length. Since Fabiani and Goulart Fabiani and Goulart (2022) considered a fixed horizon length for all case studies, no comparisons can be made between the frameworks. For Chen et al.’s Chen et al. (2022) framework, the plot clearly shows that the required network size grows even quicker with horizon length than it does with plant size. Knowing this would be particularly useful when seeking to apply the method to an existing controller. If the required horizon length is too long, the required network size could deem an approach inapplicable to the problem at hand. Lastly, Figure 5c shows how training time varies with plant size for both approximation methods. It should be noted that Fabiani and Goulart Fabiani and Goulart (2022) reported the solution times required for the MILPs involved in their approach but did not make separate reports on the training time for the DNNs. Despite this, the available data indicates that training time for their method increases faster with plant size than it does for the method by Chen et al. Chen et al. (2022). This information would be useful to know at the time of choosing an approximation framework for a new application. However, in this case the observed trends could be an artifact of the computational resources used in the respective studies. A fair comparison in training time would require performing both studies on the same hardware, and even then, differences in software implementations could skew the results.
@@
Figure 5: Examples of scalability estimates for frameworks by Chen et al. Chen et al. (2022), Fabiani and Goulart Fabiani and Goulart (2022), and Li et al. Li et al. (2022). (a) relationship between NN size and plant size. (b) relationship between NN size and prediction horizon length. (c) relationship between approximation training time and plant size. (d) relationship between training time and amount of training data, and approximation performance and amount of training data.

### v3

@@ 6.1 Methods without guarantees
As for LMPC and NMPC, the simplest way to approximate RMPC with an NN, is to use imitation learning to learn the behaviour of the controller to a satisfactory level of performance. Lucia and Karg (2018) demonstrated that this approach is applicable to RMPC despite its higher complexity. Specifically, they used a DNN with $tanh$ activation functions to learn the behaviour of multi-stage RMPC (msRMPC). Through simulation examples they proved the success of the approach at approximating the equivalent robust nonlinear controller for an industrial polymerisation reactor. The approximate controller was 2,800 times faster than the exact controller and could successfully be executed at the required rate in real time, however, it should also be pointed out that the average constraint violation, despite remaining low, was one order of magnitude larger than that of the exact controller. Thus, the pitfall of the basic imitation learning setup remains the same as previously discussed: it is not suitable for use in safety-critical systems because it provides no guarantees that the resulting approximate controller will respect the constraints of the system.
@@
In follow up work, Karg and Lucia (2019) presented a modified approach where instead of directly learning the behaviour of msRMPC, an NN was used to learn the behaviour of a slightly modified controller which contained a constraint back-off parameter that made the controller more conservative. Through simulation experiments on a small nonlinear plant with one input and three states, the authors showed that the modified imitation learning approach could be successfully used to create an approximate controller that respected the constraint to which the back-off was applied. In their example, the back-off parameter was applied to a scalar nonlinear constraint, no examples were provided for the case of multiple constraints on inputs and states. The revised method with back-off parameters is a reasonable compromise between controller conservativeness and safety. This is important because, even in systems that are not safety-critical, avoiding constraint violations may be desirable to prevent unexpected stops or mechanical failure. Nonetheless, it should be noted that the back-off strategy does not provide any formal guarantees of constraint satisfaction. By using it, constraint violation is less likely to occur and the authors outlined a method via which the probability of violation can be computed after the approximation is generated. Yet, it is important to note that the method itself does not inherently provide probabilistic or strict guarantees.
@@
Bonzanini et al. (2021) further embraced the idea that the lack of strict safety guarantees can be overlooked in favour of other valuable compromises such as enabling the use of real-time msRMPC controllers that use augmented models equipped with Gaussian Processes (GP) to learn and mitigate the effects of plant-model mismatch. The incentive for creating and using these augmented models is that the standard formulation of msRMPC assumes that the structure and distribution of uncertainties is bounded and known in advance. However, as the authors pointed out, uncertainty can be time-varying and some systems are just too hard to model using conventional approaches. In such cases GP regression can be used to learn the unmodeled dynamics or time-varying uncertainties of the system from data. The msRMPC formulation can then be modified to accommodate the GP-augmented models. Via simulation experiments on a system with significant plant-model mismatch, Bonzanini et al. showed that their learning-based reformulation of msRMPC outperformed the standard formulation but its computation cost remained too high for use in real time. To overcome this issue, they applied the basic imitation learning solution and trained a DNN to learn the behaviour of the reformulated controller. The resulting approximation enjoyed no safety guarantees, but its performance on closed-loop simulation examples matched that of the exact reformulated msRMPC controller at a quarter of the online computation cost. Karg and Lucia (2021) proposed a different method that achieves a similar result as Bonzanini et al. (2021), instead of reformulating msRMPC to account for more complex models, they trained a DNN via imitation learning to approximate the standard form of msRMPC and then used the resulting network as an initialisation of RL. Further training of the initialised agent then took care of the problem of learning the unmodeled dynamics and time-varying uncertainties. Both of these studies (Bonzanini et al., 2021; Karg and Lucia, 2021), are examples of accepting the lack of guarantees of the simplest NNBO approach to enable the use of complex controllers that build beyond classical MPC formulations.
@@ 6.2 Methods with probabilistic guarantees
In this section, we present two frameworks that are designed to preserve certain theoretical properties of RMPC on a probabilistic basis. The first framework, proposed by Hertneck et al. (2018), was an extension of the pure regression approach which includes a validation step that once passed, equips the basic approximation with statistical guarantees on constraint satisfaction and stability. In short, the validation step tests a candidate DNN approximation of the original controller on several closed-loop trajectories which start and end in feasible regions of the system’s state space, and which the original RMPC controller can solve without violating any constraints. For each trajectory in the test set, an indicator function is applied which outputs 1 if the closed-loop solution of the DNN is within a specified tolerance of the solution of the real controller, or 0 if it is not. The mean of the indicator outputs is equivalent to $\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ in HI (6). It follows that, if the value of $E(\xi)$ and the RHS of (6) are predefined, then one can calculate $\epsilon$ and verify if the difference between $\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ and $\epsilon$ is greater or equal than the prescribed $E(\xi)$ for the current number of test trajectories. If the verification is not successful, then the number of test trajectories can be increased or the quality of the DNN approximation can be improved by modifying the DNN and re-training until the validation step succeeds. If a candidate solution passes the validation step, it carries the guarantee that with some predefined confidence, its solutions are stable and constraint compliant (as defined by the indicator function) with a probability of at least the selected $E(\xi)$.
@@
Hertneck et al. (2018) applied this methodology to an RMPC formulation designed for robustness against bounded input disturbances. To preserve said robustness on the approximate controller, their choice of indicator function outputs 1 if the infinity norm of the input error between the real and approximate controllers over a closed loop trajectory is less or equal to the input disturbance bound of the original controller. Otherwise the indicator outputs 0. The effectiveness of the approach was demonstrated on a simulation study entailing the control of a continuously stirred tank reactor, in said study the performance of the approximate controller was found to be indistinguishable from that of the original policy while being over 200 times faster to evaluate.
@@
Later on, Nubert et al. (2020) applied the same approach to a more complex version of RMPC designed for setpoint tracking of nonlinear systems with robustness against plant-model mismatch. In this case the choice of indicator function was adjusted to output 1 if the solutions of the approximate controller complied with the maximum admissible model mismatch for the approximation, and 0 otherwise. The effectiveness of the reformulated approach was demonstrated on setpoint tracking experiments with a real robot arm that was represented by a simplified model with 8 states, 4 inputs and 3 outputs for control purposes. Nubert et al. (2020) reported difficulties in satisfying their HI validation criterion for high probability levels but the resulting approximate controller worked well in practice and could be executed 200 times faster than the original controller thereby allowing the use of this complex form of RMPC in real time. It should be noted that this is the only framework that has been used across multiple studies and proved to be scalable and adaptable to more complex controllers while retaining the same speed up factor. It is also one of the few frameworks that have been tested on experiments with real hardware instead of simulation.
@@
Karg et al. (2021) proposed a different framework for obtaining statistical guarantees on approximations of msRMPC which is suitable for the case where the performance of a controller with respect to a scalar indicator depends on a set of hyper-parameters. In this framework a family of controller approximations is considered, where each member of the family corresponds to a controller with a particular combination of hyper parameters. Using concepts from order statistics and probabilistic robustness analysis, the authors came up with expressions from which the size of a test set can be computed as a function of the size of the family. For said test set the authors also derived expressions to calculate the number of tests that can be failed, the desired probability of indicator violation, and the desired confidence level of the probability of violation. Each controller in the finite family can be evaluated on this test set and if the number of failed tests is lower than the allowable number of failures, then the family member is guaranteed to carry the probabilistic properties that were prescribed. Karg et al. (2021) then showed how this framework could be applied to a family of DNN approximations where each family member was trained to approximate the behaviour of an msRMPC controller with a particular constraint tightening parameter. In this example, the scalar indicator function was set to the maximum constraint violation over a 60 $s$ simulation of the closed loop system and the target performance of the probabilistic validation was to allow no violation at all with a prescribed probability. By applying their framework the size of the test set was computed and the authors showed how the different members of the family met or failed the maximum number of failures requirement. A perk of this method is that it can be applied to any arbitrary controller, not just DNN approximations of msRMPC. However, for generating neural approximations of MPC, which is the focal point of this paper, the method requires creating a data set per family member being considered. That means evaluating expensive MPC policies on large data sets multiple times instead of once like in other approximation frameworks. Therefore, there seems to be no compelling reason to choose this approximation framework over the previously discussed methods (Hertneck et al., 2018; Nubert et al., 2020) given that both can yield the same probabilistic guarantees.
@@ 6.3 Methods with strict guarantees
In this section we present methods for approximating RMPC with DNN in manners that preserve some of the theoretical properties of MPC such as the guarantees of constraint satisfaction and stability. Daosud et al. (2019) proposed an intuitive approach for preserving the stability and constraint satisfaction properties of msRMPC. Their method consists of training a DNN via imitation learning to predict the optimal cost of msRMPC for a given starting condition. They used this trained network as the terminal cost of a re-formulated msRMPC controller with a much shorter prediction horizon than the original. Shortening the prediction horizon greatly reduces the computational cost of optimisation but enables the re-formulated controller to retain the stability and constraint satisfaction properties of MPC because rigorous optimisation is still being conducted at each control interval. The authors demonstrated the effectiveness of the approach on a simulated semi-batch reactor problem consisting of a MIMO plant with dimensions as shown in Table 3. First, the original controller was implemented and simulated with a robust horizon of 1 step and a prediction horizon of 15 steps. Then, a DNN with $tanh$ activation functions was used to learn to predict the optimal cost of the solutions of this controller. Subsequently, the trained network was used as the terminal cost of a reformulated controller which had the same robust horizon as the original but a prediction horizon of only 4 steps. Through simulation experiments the authors verified that the new controller with the DNN terminal cost was stable, constraint compliant and ran 2.3 times faster on average than the original controller. The main advantages of this approach are its intuitive nature and that the computation time reductions should only improve as the length of the horizon of the original controller increases because the cost of evaluating the DNN would remain constant. The only disadvantage is that the optimality guarantees of MPC are not maintained. In their study, Daosud et al. (2019) reported a 5% performance loss when using the approximate controller compared to the original in terms of their control goal which was to maximise the concentration of a reactant in the semi-batch reactor. This performance loss was attributed to the prediction error of the DNN and it should be noted that the worst case or average performance loss were not reported, so in practice this number could be much higher, particularly if the quality of the DNN predictions worsens for harder problems and longer horizons. Nevertheless, while some performance is bound to be lost and the computation time reduction is not immense, the intuitiveness and guarantees of this method make it worthy of more investigation into its scalability and applicability to other types of MPC such as NMPC.
@@
The only other alternative with strict guarantees was proposed by Paulson and Mesbah (2020) and relies on projection operations to preserve the constraint and stability guarantees of msRMPC. Their method consists of training a DNN to learn the mapping between states and control actions for standard msRMPC. Once the network is trained its outputs are projected via optimisation onto carefully selected sets that preserve certain properties of MPC. For preserving constraint satisfaction properties in the presence of uncertainties, the authors proposed projecting the DNN outputs onto the maximal RCI set. To preserve stability properties, the authors projected the results onto the maximal $\lambda$-contractive set, which is a set in state space where all states can be driven into a tighter region of the set by applying one control input in the absence of disturbances. With this approach one can have an approximate controller that is constraint compliant in the presence of disturbances and asymptotically stable in the absence of disturbances. Nonetheless, assuming an absence of disturbances defeats the purpose of using RMPC in the first place and while the use of the $\lambda$-contractive set can yield Input-to-State Stability (ISS) guarantees for disturbed cases, there are additional conditions that need to be satisfied in this case which are difficult and perhaps impossible to verify a priori. In fact, the requirements for the constraint satisfaction guarantees are also hard to meet because the RCI set cannot always be computed or approximated for complex system with large plants. The authors demonstrated their approach on a double integrator plant problem and reported a computation speed up of 400 times compared to the original controller when the optimisation based projection operation was solved offline as an mpQP. No speed up factors or results were reported for cases where the projection operation cannot be solved offline. The applicability of this method to more complex problems has not been studied but seems unlikely.
@@ 7.1 Problem types and the lack of benchmarks
From the graph it is clear that there are a few problem types that are significantly more popular, while other types have only been encountered once or twice. The trend we aim to highlight is the absence of well-established benchmarks for testing approximation frameworks. The stirred or semi-batch tank reactor problems are by far the most commonly encountered problem type in case studies. Ren et al. (2022) made the same observation on their survey on NN modelling approaches for MPC. Specifically, they noted that chemical reaction problems are commonly encountered in case studies of machine learning-assisted optimal control because it is difficult to model such processes from first principles. In the works covered in this review, versions of the stirred reactor problem type are usually polymerisation reactions or pH control problems. Although the plants involved vary in complexity, most are nonlinear MIMO plants of medium size with 4 to 8 states and 2 to 3 inputs. The second most popular problem type is the double integrator, which is a SISO plant with 2 states. This problem type is not dissimilar from undamped oscillator problems. We encountered some publications (Paulson and Mesbah, 2020; Karg and Lucia, 2020b; Pin et al., 2010, 2013) that demonstrated their approach only on these problem types, which raises concerns about their scalability. Heating system problems are the third most popular type. The plants involved in these problems vary in size, ranging from medium to very large with hundreds of states. It is evident then that there is no standardised benchmark for showcasing the effectiveness of a given framework.
@@
Some research studies addressed the lack of benchmarks challenge by using multiple examples and more than one problem type to demonstrate an approach. For instance, Fabiani and Goulart (2022) used six versions of the oscillating masses problem with increasing dimensions and a fixed horizon length to demonstrate the scalability of their method. Similarly, Chen et al. (2022) considered three problem types and four case studies including the double integrator, quadrotor, and two instances of the oscillating masses problem. By demonstrating their approach on plants of different sizes and different nature, they not only proved the scalability of their method, but also its applicability beyond a specific problem type. Alternatively, Li et al. (2022) focused on a single plant but made efforts to demonstrate the scalability of their method in different ways. Most notably, they showed how the training time and performance of their approach varied as more training data was considered. Studies such as these allow to make important comparisons between approximation methods. In turn, these comparisons shed light on the information that a comprehensive benchmark should aim to measure.
@@
Figure 5 shows several relationships estimated using the data extracted from the case studies by Fabiani and Goulart (2022); Chen et al. (2022) and Li et al. (2022). Figures 5a – c compare the frameworks by Chen et al. (2022) and Fabiani and Goulart (2022) which are both of the strict guarantees type for LMPC, but are otherwise very different in nature. The former consists of training a DNN that predicts all primal variables to hot-start an optimisation solver. The latter proposes MILPs to compute the values of properties that can be used to verify the stability and constraint compliance of controllers approximated via DNN ReLU regression. As shown in the plots, these differences in nature have profound impacts on the way that both methods scale for problems of different size. For example, the equations of the fitted functions in Figure 5a show that the required network size in the method by Chen et al. (2022) scales exponentially with plant size. However, for the method by Fabiani and Goulart (2022) it scales linearly. This information is valuable because training large networks is likely to require more time and computational resources than for smaller networks. Note however, that because Fabiani and Goulart (2022) only considered one problem type of various sizes, it is not possible to say if this linear scaling holds in general. Figure 5b shows how the required NN size varies with prediction horizon length. Since Fabiani and Goulart (2022) considered a fixed horizon length for all case studies, no comparisons can be made between the frameworks. For the Chen et al. (2022) framework, the plot clearly shows that the required network size grows even quicker with horizon length than it does with plant size. Knowing this would be particularly useful when seeking to apply the method to an existing controller. If the required horizon length is too long, the required network size could deem an approach inapplicable to the problem at hand. Lastly, Figure 5c shows how training time varies with plant size for both approximation methods. It should be noted that Fabiani and Goulart (2022) reported the solution times required for the MILPs involved in their approach but did not make separate reports on the training time for the DNNs. Despite this, the available data indicates that training time for their method increases faster with plant size than it does for the method by Chen et al. (2022). This information would be useful to know at the time of choosing an approximation framework for a new application. However, in this case the observed trends could be an artifact of the computational resources used in the respective studies. A fair comparison in training time would require performing both studies on the same hardware, and even then, differences in software implementations could skew the results.
@@
Figure 5: Examples of scalability estimates for frameworks by Chen et al. (2022), Fabiani and Goulart (2022), and Li et al. (2022). (a) relationship between NN size and plant size. (b) relationship between NN size and prediction horizon length. (c) relationship between approximation training time and plant size. (d) relationship between training time and amount of training data, and approximation performance and amount of training data.